The longest mounted ladder used by the Dallas Fire Department is 108 ft long. The optimum operating angle of this ladder is 60 degrees. The ladder is mounted on the back of the 8 ft long truck.
A. What is the tallest building which the 108 ft ladder can reach using the optimum operating angle?
B. How far away from the building would the ladder be placed?

Answer by mananth(12270)

The longest mounted ladder used by the Dallas Fire Department is 108 ft long. The optimum operating angle of this ladder is 60 degrees. The ladder is mounted on the back of the 8 ft long truck.
A. What is the tallest building which the 108 ft ladder can reach using the optimum operating angle?
B. How far away from the building would the ladder be placed?
ladder length =108
optimum angle = 60 deg.
vertical height = 108* cos 60
= 108*0.866
= 93.5 feet is the maximum height of the building it can approach
maximum height = 93.5
length of ladder =108 (hypo +8tenuse)

tan 60 = vertical height / base
base = vertical height / tan 60
base = 93.5 / 1.7321
= 53.99 feet - distance friom the building to foot of the ladder